# A Bayesian Modeling Framework for Estimation and Ground Segmentation of Cluttered Staircases

### v1

@@ Abstract
Autonomous robot navigation in complex environments requires robust perception as well as high-level scene understanding due to perceptual challenges, such as occlusions, and uncertainty introduced by robot movement. For example, a robot climbing a cluttered staircase can misinterpret clutter as a step, misrepresenting the state and compromising safety. This requires robust state estimation methods capable of inferring the underlying structure of the environment even from incomplete sensor data. In this paper, we introduce a novel method for robust state estimation of staircases. To address the challenge of perceiving occluded staircases extending beyond the robot’s field-of-view, our approach combines an infinite-width staircase representation with a finite endpoint state to capture the overall staircase structure. This representation is integrated into a Bayesian inference framework to fuse noisy measurements enabling accurate estimation of staircase location even with partial observations and occlusions. Additionally, we present a segmentation algorithm that works in conjunction with the staircase estimation pipeline to accurately identify clutter-free regions on a staircase. Our method is extensively evaluated on real robot across diverse staircases, demonstrating significant improvements in estimation accuracy and segmentation performance compared to baseline approaches.
@@
Staircase Modeling, Staircase Estimation, Legged Robots, Probabilistic Robot Perception, Ground Segmentation
@@ I Introduction
Figure 1: (a) Prior staircase estimate, $bel(\boldsymbol{{}_{L}X}_{k})$ at time $k$ (solid blue) (b) Predicted staircase location, $p(\boldsymbol{{}_{L}\hat{X}}_{k+1|k})$ based on prior (dotted blue) (c) New noisy measurement, $p(\boldsymbol{{}_{L}Z}_{k+1})$ at time $k\!+\!1$ (magenta) (d) Filtered staircase estimate $bel(\boldsymbol{{}_{L}X}_{k+1})$ at time $k\!+\!1$ by combing prediction and measurement (green)
Figure 2: Results of our proposed staircase modeling and estimation framework (a) Robot navigating staircase environments (b) Estimated staircase highlighted by white marker (c) Segmented clutter-free regions on each stair
@@
We implement our proposed method both on real robots and in simulation in diverse environments with large and cluttered staircases (Fig. 2). We achieve real-time performance with a 30-94% reduction in estimated staircase parameter and location errors compared to baselines.
@@ II-A Staircase Detection and Estimation
Researchers have explored various methods for staircase detection using both image-based and point cloud-based approaches. While image-based methods [Murakami] [ilyas2023staircase] offer computational speed, they are susceptible to environmental factors like lighting conditions, and face challenges in accurately estimating 3D geometry. Point cloud methods leverage the inherent planar geometry of stairs to segment planes and estimate staircase location and parameters. Although RANSAC-based approaches are prevalent [sanchez2021staircase] [fourre2020autonomous], they can be non-deterministic and computationally expensive. Westfechtel et al. [westfechtel2018robust] achieved high accuracy in staircase parameter estimation using a graph-based search strategy, but required significant processing time. Qing et al. [qing2024onboard] introduced a measurement correction technique based on estimated stair plane poses, but their approach is sensitive to plane segmentation errors that can be caused by clutter on stairs.
The authors’ prior work in [sriganesh2023fast] presented a fast staircase detection method and introduced a simple matching algorithm to combine multiple detections over time, which averaged individual matched stairs based on proximity. However, this approach proved susceptible to noise, leading to estimation failures due to erroneous matches. Furthermore, the absence of explicit staircase modeling and the inability to leverage existing staircase information limited its capacity to handle incomplete detections arising from the sensor’s restricted field of view or occlusions.
@@ II-B Ground Segmentation
Extensive research has investigated ground segmentation techniques due to their critical role in ensuring safe robot navigation [gomes2023survey]. Elevation maps are a simple and popular approach, which transform 3D point clouds into 2.5D grid representations for ground point classification [elevation_map1]. Plane fitting techniques, frequently employing RANSAC, classify points based on their proximity to a fitted plane [plane_fitting1]. These methods encounter difficulties in scenarios with multiple ground planes, such as staircases, as they necessitate multiple plane estimations, which is computationally expensive. While variations like concentric zone divisions [lim2021patchwork] strive to accelerate this process, the inherent complexity of cluttered staircases persists as a challenge.
More sophisticated methods, including Markov Random Fields (MRFs) [huang2021fastmrf] and Gaussian processes [chen2014gaussian], offer enhanced accuracy but demand substantial computational resources, posing challenges for mobile robots with limited compute. Cloth simulation filters [zhang2016csf] present an alternative, yet their performance on stairs is highly dependent on parameter tuning. Furthermore, purely geometric methods struggle to differentiate between small, flat obstacles on stairs and adjacent steps, raising safety concerns. Learning-based methods [milioto2019rangenet++][he2022sectorgsnet] integrate semantic information to improve performance, but their effectiveness remains contingent on the specific scenarios encountered during training.
@@ III-A Staircase State-Space
(a) Top down view
(b) 3-D view
Figure 3: Staircase annotated with the ‘infinite-line’ and ‘staircase endpoint’ state variables
We present a split state-space model for representing staircases, enabling a robot to capture and reason about those extending beyond its sensor range. The first component, the ‘infinite-line staircase state’ ($\mathbf{{}_{L}X}$) , represents the intersection of each stair’s tread and riser as an infinite line using polar coordinates. Subsequently, for each stair $i$ in the staircase, its state $\boldsymbol{{\bar{\xi}}}^{(i)}$ consists of radius $r^{(i)}$, its angle $\phi^{(i)}$, along with the start and end z-positions $(z_{s}^{(i)},z_{e}^{(i)})$ of each stair, all with respect to the robot’s world frame. Fig. 3(a) illustrates this infinite-line state for a small staircase. For a staircase with $N$ steps at time step $k$, the infinite-line state vector, $\mathbf{{}_{L}X}_{k}$, consists of each stair line in ascending staircase direction.
$\mathbf{{}_{L}X}_{k}=\begin{bmatrix}\boldsymbol{\bar{\xi}}^{(1)}_{k}\;\dots\;\boldsymbol{\bar{\xi}}^{(N)}_{k}\end{bmatrix},\boldsymbol{\bar{\xi}}^{(i)}_{k}=\begin{array}[]{c}[r^{(i)}\;\;\phi^{(i)}\;\;z^{(i)}_{s}\;\;z_{e}^{(i)}]\end{array}$ (2)
The second component, the ‘staircase endpoint state’ ($\mathbf{{}_{P}X}$), grounds this representation by incorporating the physical location of each step. $\mathbf{{}_{P}X}$ uses cartesian coordinates to define the actual starting ($\mathbf{\bar{p}_{s}}$) and ending points ($\mathbf{\bar{p}_{e}}$) of each stair line with respect to the robot’s world frame. This precise 3D information allows for accurate reconstruction of parts of the staircase that has been observed, while the infinite-line state can help describe potentially unseen parts of the staircase. Fig. 3 illustrates this staircase endpoint state for a small staircase. For a staircase with $N$ steps at time step $k$, the endpoint state vector, $\mathbf{{}_{P}X}_{k}$, consists of each stair’s endpoints in ascending staircase direction.
$\mathbf{{}_{P}X}_{k}$ $=\begin{bmatrix}\mathbf{\bar{p}}_{s}^{(1)}&\mathbf{\bar{p}}_{e}^{(1)}&\dots&\mathbf{\bar{p}}_{s}^{(N)}&\mathbf{\bar{p}}_{e}^{(N)}\end{bmatrix},$ (3)
$\mathbf{\bar{p}}_{s}^{(i)}$ $=\begin{array}[]{c}[\;x_{s}^{(i)}\;\;y_{s}^{(i)}\;\;z^{(i)}_{s}\;]\end{array},\;\;\;\mathbf{\bar{p}}_{e}^{(i)}=\begin{array}[]{c}[\;x_{e}^{(i)}\;\;y_{e}^{(i)}\;\;z_{e}^{(i)}\;]\end{array}$ (6)
@@ III-B Parameterized Staircase Model
We model a staircase using six key parameters $\mathbf{S}_{k}$. We assume that the parameters remain consistent between any two consecutive steps along the staircase. As depicted in Fig. 4, the six parameters of our staircase model are:
@@
  • Stair Depth ($d$): The horizontal depth of each step. In the case of curved staircases, the depth is measured from the center of each stair line, as illustrated in Fig. 4. This ensures consistent measurement even when the stair edges are not parallel.
@@
Each of these parameters can be computed directly from the current staircase state $(\mathbf{{}_{L}X}_{k},\mathbf{{}_{P}X}_{k})$. To minimize the impact of potential outliers, the stair height, depth, width, and curvature angle of the staircase are determined by averaging the corresponding values across all individual steps. It’s important to distinguish the stair start and end angles $(\psi_{s},\psi_{e})$ from the infinite-line state angles $(\phi^{(1)},\phi^{(N)})$. While $\psi_{s}$ and $\psi_{e}$ represent the orientation of the perpendicular to the stair line in the ascending direction of the staircase, $\phi^{(i)}$ simply represents the angle of the normal to the line. Hence, based on the staircase’s location in the map $\psi_{s}$ can be the same as $\phi^{(1)}$ or $\phi^{(1)}+\pi$. The same applies to $\psi_{e}$ as well.
This parameterized model is crucial for robustly estimating individual stair locations, as it allows us to predict the location of a stair using information from its immediate neighbors. $\boldsymbol{\bar{\xi}}^{(i,j)}$ denotes the state of stair $i$ predicted given the state of its adjacent stair $j$ using the stair parameters, as shown in (10). These equations incorporate several parameters to generalize for diverse staircase configurations. The parameter $\eta$ controls the direction of prediction, where $\eta=1$ for predicting a succeeding stair ($j=i-1$) and $\eta=-1$ for a preceding stair $(j=i+1)$. As detailed before, the orientation of stair $j$ (given by $\psi_{s}\pm j\Delta\psi$) can differ from the actual angle of the stair line ($\phi^{(j)}$) by $\pi$, depending on the staircase’s position and the world frame’s orientation. To account for this, the parameter $\rho$ is introduced. Moreover, since a curved staircase has different depths at different points on the line, the parameters $\gamma$ and $l$ account for this difference and correct the stair depth ($d$) based on distance from the center of the given stair ($c_{x}^{(j)},c_{y}^{(j)}$).
@@
Similarly, we can predict the endpoint state for stair $i$ given neighboring stair $j$, denoted as $\mathbf{\bar{p}}_{s}^{(i,j)}$ and $\mathbf{\bar{p}}_{e}^{(i,j)}$ as shown in (11). The prediction starts from the known endpoint of stair $j$ and extrapolates them based on the direction. The predictions conform to the direction of extrapolation, as it utilizes the starting stair yaw angle ($\psi_{s}$) when predicting a succeeding stair and the final stair yaw angle ($\psi_{e}$) when predicting a preceding stair. Furthermore, the prediction incorporates variations in step depth ($d$) for curved staircases by accounting for the curvature ($\Delta\psi$).
Figure 4: Illustration depicting all the parameters for a staircase
@@
By iteratively applying these equations, these equations can be extended to estimate the state of non-adjacent stairs (i.e., stairs where $j$ is not an immediate neighbor of $i$), effectively enabling the prediction of any stair along the staircase. This is particularly useful when needing to predict the state of multiple new stairs based on the information from a single observed stair.
@@ III-C Staircase Measurements and Initialization
To enable state estimation, we leverage staircase detections from our staircase detection algorithm [sriganesh2023fast] as measurements. Mirroring our state representation, these measurements are split into two components: measured stair lines ($\mathbf{{}_{L}Z}_{k}$) and measured stair endpoints ($\mathbf{{}_{P}Z}_{k}$). However, unlike the state representation, these measurements are represented in the robot’s local frame. This aids in decoupling uncertainty in detection and the robot localization uncertainty. The robot’s local frame is defined with its x-y plane parallel to the world x-y plane, simplifying the corresponding robot’s pose ($\boldsymbol{\Omega}_{k}$) to its position ($\bar{t}_{r}$) and yaw orientation ($\theta_{r}$).
@@
$\mathbf{{}_{P}Z}_{k}$ $=\begin{bmatrix}\mathbf{\bar{p}}_{sm}^{(1)}&\mathbf{\bar{p}}_{em}^{(1)}&\dots&\mathbf{\bar{p}}_{sm}^{(M)}&\mathbf{\bar{p}}_{em}^{(M)}\end{bmatrix},$ (14)
$\mathbf{\bar{p}}_{sm}^{(i)}$ $=\begin{array}[]{c}[\;x_{sm}^{(i)}\;\;y_{sm}^{(i)}\;\;z^{(i)}_{sm}\;]\end{array},\;\;\mathbf{\bar{p}}_{em}^{(i)}=\begin{array}[]{c}[\;x_{em}^{(i)}\;\;y_{em}^{(i)}\;\;z_{em}^{(i)}\;]\end{array}\!,$ (17)
$\boldsymbol{\Omega}_{k}$ $=\begin{bmatrix}\bar{t}_{r}&\theta_{r}\end{bmatrix}=\begin{bmatrix}x_{r}&y_{r}&z_{r}&\theta_{r}\end{bmatrix}$ (18)
We represent the belief of the infinite line state ($bel(\boldsymbol{{}_{L}X}_{k})$) at time $k$ by the mean $\boldsymbol{{}_{L}X}_{k}$ and the covariance $\boldsymbol{{}_{L}\Sigma}_{k}$. Upon initial detection of a staircase, we initialize the infinite-line staircase state ($\mathbf{{}_{L}X}_{k}$) using the robot pose ($\boldsymbol{\Omega}_{k}$) as described in (19). We also initialize a state covariance matrix ($\mathbf{{}_{L}\Sigma}_{k}$) for the infinite-line state as shown in (20). This covariance incorporates both the inherent noise in the measurements ($\mathbf{\mathcal{Q}}$) and the uncertainty in the estimated robot pose from SLAM ($\mathbf{\Sigma}_{\boldsymbol{\Omega}}$). We model the measurement noise as zero-mean Gaussian with standard deviations $\sigma_{r}$, $\sigma_{\phi}$, $\sigma_{z_{s}}$, $\sigma_{z_{e}}$ for each component of the measured stair line ($\mathbf{{}_{L}Z}_{k}$). We also initialize the endpoint state ($\mathbf{{}_{P}X}_{k}$) as shown in (21).
$\boldsymbol{\bar{\xi}}^{(i)}_{k}$ $=g(\boldsymbol{\zeta}^{(i)}_{k},\boldsymbol{\Omega}_{k})$ (19)
@@
$\mathbf{\bar{p}}_{s}^{(i)}$ $=\mathbf{R}(\theta_{r})\mathbf{\bar{p}}_{sm}^{(i)}+\bar{t_{r}},\;\;\;\;\mathbf{\bar{p}}_{e}^{(i)}=\mathbf{R}(\theta_{r})\mathbf{\bar{p}}_{em}^{(i)}+\bar{t_{r}}$ (21)
$\mathbf{R}(\theta_{r})\textit{ is the SO3 rotation matrix with yaw }\theta_{r}$
It is important to note that, unlike the infinite-line state, we do not maintain a covariance matrix for the stair endpoint state. This design choice is motivated by two factors. Firstly, the weighted line fitting algorithm [pfister2003weighted], employed by our staircase detection module [sriganesh2023fast], inherently handles the noise in polar coordinates during the segmentation process. Secondly, we prioritize refining the infinite-line state via the Extended Kalman Filter (EKF) (explained in Section IV) and subsequently update the endpoints based on this refined state.

### v2

@@ Abstract
Autonomous robot navigation in complex environments requires robust perception as well as high-level scene understanding due to perceptual challenges, such as occlusions, and uncertainty introduced by robot movement. For example, a robot climbing a cluttered staircase can misinterpret clutter as a step, misrepresenting the state and compromising safety. This requires robust state estimation methods capable of inferring the underlying structure of the environment even from incomplete sensor data. In this paper, we introduce a novel method for robust state estimation of staircases. To address the challenge of perceiving occluded staircases extending beyond the robot’s field-of-view, our approach combines an infinite-width staircase representation with a finite endpoint state to capture the overall staircase structure. This representation is integrated into a Bayesian inference framework to fuse noisy measurements enabling accurate estimation of staircase location even with partial observations and occlusions. Additionally, we present a segmentation algorithm that works in conjunction with the staircase estimation pipeline to accurately identify clutter-free regions on a staircase. Our method is extensively evaluated on real robots across diverse staircases, demonstrating significant improvements in estimation accuracy and segmentation performance compared to baseline approaches.
@@
Object Detection, Segmentation and Categorization; Probabilistic Inference; Field Robots
@@ I Introduction
Figure 1: (a) Prior staircase estimate, $bel(\boldsymbol{{}_{L}X}_{k})$ at time $k$ (solid blue). (b) Predicted staircase location, $p(\boldsymbol{{}_{L}\hat{X}}_{k+1|k})$ based on prior (dotted blue). (c) New noisy measurement, $p(\boldsymbol{{}_{L}Z}_{k+1})$ at time $k\!+\!1$ (magenta). (d) Filtered staircase estimate $bel(\boldsymbol{{}_{L}X}_{k+1})$ at time $k\!+\!1$ by combing prediction and measurement (green).
Figure 2: Results of our proposed staircase modeling and estimation framework. (a) Robot navigating staircase environments (b) Estimated staircase highlighted by white marker. (c) Segmented clutter-free regions on each stair.
@@
We implement our proposed method both on real robots and in simulation in diverse environments with large and cluttered staircases (Fig. 2). We achieve real-time performance with 67-89% reduction in estimated staircase parameter errors and 30% reduction in staircase location error compared to baselines
@@ II-A Staircase Detection and Estimation
Researchers have explored various methods for staircase detection using both image-based and point cloud-based approaches. While image-based methods [1] [2] offer computational speed, they are susceptible to environmental factors like lighting conditions, and face challenges in accurately estimating 3D geometry. Point cloud methods leverage the inherent planar geometry of stairs to segment planes and estimate staircase location and parameters. Although RANSAC-based approaches are prevalent [3] [4], they can be non-deterministic and computationally expensive. Perez-Yus et al. used a wearable RGB-D camera to detect staircases [5]. Westfechtel et al. [6] achieved high accuracy in staircase parameter estimation using a graph-based search strategy, but required significant processing time. Qing et al. [7] introduced a measurement correction technique based on estimated stair plane poses, but their approach is sensitive to plane segmentation errors that can be caused by clutter on stairs.
The authors’ prior work in [8] presented a fast staircase detection method and introduced a simple matching algorithm to combine multiple detections over time, which averaged individual matched stairs based on proximity. However, this approach proved susceptible to noise, leading to estimation failures due to erroneous matches. Furthermore, the absence of explicit staircase modeling and the inability to leverage existing staircase information limited its capacity to handle incomplete detections arising from the sensor’s restricted field of view or occlusions.
@@ II-B Ground Segmentation
Extensive research has investigated ground segmentation techniques due to their critical role in ensuring safe robot navigation [9]. Elevation maps are a simple and popular approach, which transform 3D point clouds into 2.5D grid representations for ground point classification [10]. Plane fitting techniques, frequently employing RANSAC, classify points based on their proximity to a fitted plane [11]. These methods encounter difficulties in scenarios with multiple ground planes, such as staircases, as they necessitate multiple plane estimations, which is computationally expensive. While variations like concentric zone divisions [12] strive to accelerate this process, the inherent complexity of cluttered staircases persists as a challenge.
More sophisticated methods, including Markov Random Fields (MRFs) [13] and Gaussian processes [14], offer enhanced accuracy but demand substantial computational resources, posing challenges for mobile robots with limited compute. Cloth simulation filters [15] present an alternative, yet their performance on stairs is highly dependent on parameter tuning. Furthermore, purely geometric methods struggle to differentiate between small, flat obstacles on stairs and adjacent steps, raising safety concerns. Learning-based methods [16][17] integrate semantic information to improve performance, but their effectiveness remains contingent on the specific scenarios encountered during training.
@@ III-A Staircase State-Space
(a) Top down view of the staircase.
(b) 3-D view of the staircase.
(c) Model parameters for a staircase.
Figure 3: Staircase annotated with the ‘infinite-line’ and ‘staircase endpoint’ state variables along with its model parameters.
We present a split state-space model for representing staircases, enabling a robot to capture and reason about those extending beyond its sensor range. The first component, the ‘infinite-line staircase state’ ($\mathbf{{}_{L}X}$) , represents the intersection of each stair’s tread and riser as an infinite line using polar coordinates. Subsequently, for each stair $i$ in the staircase, its state $\boldsymbol{\xi}^{(i)}$ consists of radius $r^{(i)}$, its angle $\phi^{(i)}$, along with the start and end z-positions $(z_{s}^{(i)},z_{e}^{(i)})$ of each stair, all with respect to the robot’s world frame. The infinite-line state is illustrated in Fig. 3(a). For a staircase with $N$ steps at time step $k$, the infinite-line state vector, $\mathbf{{}_{L}X}_{k}$, consists of each stair line in ascending staircase direction.
$\mathbf{{}_{L}X}_{k}=\begin{bmatrix}\boldsymbol{\xi}^{(1)}_{k}\;\dots\;\boldsymbol{\xi}^{(N)}_{k}\end{bmatrix},\boldsymbol{\xi}^{(i)}_{k}=\begin{array}[]{c}[r^{(i)}\;\;\phi^{(i)}\;\;z^{(i)}_{s}\;\;z_{e}^{(i)}]\end{array}$ (2)
The second component, the ‘staircase endpoint state’ ($\mathbf{{}_{P}X}$), grounds this representation by incorporating the physical location of each step. $\mathbf{{}_{P}X}$ uses cartesian coordinates to define the actual starting ($\mathbf{p}_{s}$) and ending points ($\mathbf{p}_{e}$) of each stair line with respect to the robot’s world frame (Fig. 3). This precise 3D information allows for accurate reconstruction of parts of the staircase that have been observed, while the infinite-line state can help describe potentially unseen parts of the staircase. For a staircase with $N$ steps at time step $k$, the endpoint state vector, $\mathbf{{}_{P}X}_{k}$, consists of each stair’s endpoints in ascending staircase direction.
$\mathbf{{}_{P}X}_{k}$ $=\begin{bmatrix}\mathbf{p}_{s}^{(1)}&\mathbf{p}_{e}^{(1)}&\dots&\mathbf{p}_{s}^{(N)}&\mathbf{p}_{e}^{(N)}\end{bmatrix},$ (3)
$\mathbf{p}_{s}^{(i)}$ $=\begin{array}[]{c}[\;x_{s}^{(i)}\;\;y_{s}^{(i)}\;\;z^{(i)}_{s}\;]\end{array},\;\;\;\mathbf{p}_{e}^{(i)}=\begin{array}[]{c}[\;x_{e}^{(i)}\;\;y_{e}^{(i)}\;\;z_{e}^{(i)}\;]\end{array}$ (6)
As our model represents stairs as edges (lines) it excludes the information of the initial ground plane for ascending staircases. Similarly, for descending staircases, it excludes the final ground plane at the bottom.
@@ III-B Parameterized Staircase Model
We model a staircase using six key parameters $\mathbf{S}_{k}$. We assume that the parameters remain consistent between any two consecutive steps along the staircase. As depicted in Fig. 3(c), the six parameters of our staircase model are:
@@
  • Stair Depth ($d$): The horizontal depth of each step. In the case of curved staircases, the depth is measured from the center of each stair line, as illustrated in Fig. 3(c). This ensures consistent measurement even when the stair edges are not parallel.
@@
Each of these parameters can be computed directly from the current staircase state $(\mathbf{{}_{L}X}_{k},\mathbf{{}_{P}X}_{k})$. To minimize the impact of noise, the stair height, depth, width, and curvature angle of the staircase are determined by averaging the corresponding values across all individual steps. Staircases divided by a landing are treated as multiple, separate staircases. The pronounced difference in stair depth at a landing is used to detect these instances. It’s important to distinguish the stair start and end angles $(\psi_{s},\psi_{e})$ from the infinite-line state angles $(\phi^{(1)},\phi^{(N)})$. While $\psi_{s}$ and $\psi_{e}$ represent the orientation of the perpendicular to the stair line in the ascending direction of the staircase, $\phi^{(i)}$ simply represents the angle of the normal to the line. Hence, based on the staircase’s location in the map $\psi_{s}$ can be the same as $\phi^{(1)}$ or $\phi^{(1)}+\pi$. The same applies to $\psi_{e}$ as well.
This parameterized model is crucial for robustly estimating individual stair locations, as it allows us to predict the location of a stair using information from its immediate neighbors. $\boldsymbol{\xi}^{(i,j)}$ denotes the state of stair $i$ predicted given the state of its adjacent stair $j$ using the stair parameters, as shown in (10). These equations incorporate several parameters to generalize for diverse staircase configurations. The parameter $\eta$ controls the direction of prediction, where $\eta=1$ for predicting a succeeding stair ($j=i-1$) and $\eta=-1$ for a preceding stair $(j=i+1)$. As detailed before, the orientation of stair $j$ (given by $\psi_{s}\pm j\Delta\psi$) can differ from the actual angle of the stair line ($\phi^{(j)}$) by $\pi$, depending on the staircase’s position and the world frame’s orientation. To account for this, the parameter $\rho$ is introduced. Moreover, since a curved staircase has different depths at different points on the line, the parameters $\gamma$ and $l$ account for this difference and correct the stair depth ($d$) based on distance from the center of the given stair ($c_{x}^{(j)},c_{y}^{(j)}$).
@@
Similarly, we can predict the endpoint state for stair $i$ given neighboring stair $j$, denoted as $\mathbf{p}_{s}^{(i,j)}$ and $\mathbf{p}_{e}^{(i,j)}$ as shown in (11). The prediction starts from the known endpoint of stair $j$ and extrapolates them based on the direction. The predictions conform to the direction of extrapolation, as it utilizes the starting stair yaw angle ($\psi_{s}$) when predicting a succeeding stair and the final stair yaw angle ($\psi_{e}$) when predicting a preceding stair. Furthermore, the prediction incorporates variations in step depth ($d$) for curved staircases by accounting for the curvature ($\Delta\psi$).
@@
We can estimate the state of non-adjacent stairs (i.e., stairs where $j$ is not an immediate neighbor of $i$) by iteratively applying (11), effectively enabling the prediction of any stair along the staircase. This is particularly useful to predict the state of multiple new stairs based on the information from a single observed stair.
@@ III-C Staircase Measurements and Initialization
We leverage raw detections from our staircase detection algorithm [8] as measurements ($\mathbf{{}_{L}Z}_{k}$, $\mathbf{{}_{P}Z}_{k}$). Measurements only capture the portion of the staircase within the robot’s current field-of-view and are expressed in the robot’s local frame. They are split into two components: measured stair lines ($\mathbf{{}_{L}Z}_{k}$) and measured stair endpoints ($\mathbf{{}_{P}Z}_{k}$). In contrast, the staircase state $(\mathbf{{}_{L}X},\mathbf{{}_{P}X})$ provides an estimate of the entire staircase structure in the world coordinate frame. This local frame measurement approach aids in decoupling uncertainty in detection and the robot localization uncertainty. The robot’s local frame is defined with its x-y plane parallel to the world x-y plane, simplifying the corresponding robot’s pose ($\boldsymbol{\Omega}_{k}$) to its position ($t_{r}$) and yaw orientation ($\theta_{r}$).
@@
$\mathbf{{}_{P}Z}_{k}$ $=\begin{bmatrix}\mathbf{p}_{sm}^{(1)}&\mathbf{p}_{em}^{(1)}&\dots&\mathbf{p}_{sm}^{(M)}&\mathbf{p}_{em}^{(M)}\end{bmatrix},$ (14)
$\mathbf{p}_{sm}^{(i)}$ $=\begin{array}[]{c}[\;x_{sm}^{(i)}\;\;y_{sm}^{(i)}\;\;z^{(i)}_{sm}\;]\end{array},\;\;\mathbf{p}_{em}^{(i)}=\begin{array}[]{c}[\;x_{em}^{(i)}\;\;y_{em}^{(i)}\;\;z_{em}^{(i)}\;]\end{array}\!,$ (17)
$\boldsymbol{\Omega}_{k}$ $=\begin{bmatrix}t_{r}&\theta_{r}\end{bmatrix}=\begin{bmatrix}x_{r}&y_{r}&z_{r}&\theta_{r}\end{bmatrix}$ (18)
To enable state estimation, we represent the infinite line state by its belief, $bel(\boldsymbol{{}_{L}X}_{k})$ at time $k$ with mean $\boldsymbol{{}_{L}X}_{k}$ and the covariance $\boldsymbol{{}_{L}\Sigma}_{k}$. Upon initial detection of a staircase, we initialize the infinite-line staircase state ($\mathbf{{}_{L}X}_{k}$) by transforming the measurement to world frame. This transformation, $g(\boldsymbol{\zeta}^{(i)}_{k},\boldsymbol{\Omega}_{k})$ is shown in (19). We also initialize a state covariance matrix ($\mathbf{{}_{L}\Sigma}_{k}$) for the infinite-line state as shown in (20). This covariance incorporates both the inherent noise in the measurements ($\mathbf{\mathcal{Q}}$) and the uncertainty in the estimated robot pose from SLAM ($\mathbf{\Sigma}_{\boldsymbol{\Omega}}$). We model the measurement noise as zero-mean Gaussian with standard deviations $\sigma_{r}$, $\sigma_{\phi}$, $\sigma_{z_{s}}$, $\sigma_{z_{e}}$ for each component of the measured stair line ($\mathbf{{}_{L}Z}_{k}$). We also initialize the endpoint state ($\mathbf{{}_{P}X}_{k}$) as shown in (21).
$\boldsymbol{\xi}^{(i)}_{k}$ $=g(\boldsymbol{\zeta}^{(i)}_{k},\boldsymbol{\Omega}_{k})$ (19)
@@
$\mathbf{p}_{s}^{(i)}$ $=\mathbf{R}(\theta_{r})\mathbf{p}_{sm}^{(i)}+t_{r},\;\;\;\;\mathbf{p}_{e}^{(i)}=\mathbf{R}(\theta_{r})\mathbf{p}_{em}^{(i)}+t_{r}$ (21)
$\mathbf{R}(\theta_{r})\textit{ is the SO3 rotation matrix with pure yaw }\theta_{r}$
It is important to note that, unlike the infinite-line state, we do not maintain a covariance matrix for the stair endpoint state. This design choice is motivated by two factors. Firstly, the weighted line fitting algorithm [18], employed by our staircase detection module [8], inherently handles the noise in polar coordinates during the segmentation process. Secondly, we prioritize refining the infinite-line state via the Extended Kalman Filter (EKF) (explained in Section IV) and subsequently update the endpoints based on this refined state.